A sample of helium gas has a temperature of 298 K and a pressure of 3.60 atm, under a constant volume. What is the new temperature, if the pressure is increased to 3.80 atm?
Answer #1

PV = nRT

At constant volume, P = T

P1/P2 = T1/T2

3.60 atm / 3.80 atm = 298 K / T2

0.9474 = 298 / T2

T2 = 298 / 0.9474 = 314.5 K

Therefore, new temperature = 314.5 K